How long will it take B dollars to double itself at 15% simple interest?
Answer by mananth(16949)

How long will it take B dollars to double itself at 15% simple interest?
Principal = b
Interest = b ( since the principal doubles)
N = number of years = 100 * SI / P*R (SI = simple interest R= rate of interest)

100 *b / b*15
= 100/15
= 6.67 years
enjoy your maths